# Optimistic Thompson Sampling for No-Regret Learning in Unknown Games

### 07 Feb 2024 v1

@@ Abstract
Many real-world problems involving multiple decision-makers can be modeled as an unknown game characterized by partial observations. Addressing the challenges posed by partial information and the curse of multi-agency, we developed Thompson sampling-type algorithms, leveraging information about opponent’s action and reward structures. Our approach significantly reduces experimental budgets, achieving a more than tenfold reduction compared to baseline algorithms in practical applications like traffic routing and radar sensing. We demonstrate that, under certain assumptions about the reward structure, the regret bound exhibits merely a logarithmic dependence on the total action space size, effectively mitigating the curse of multi-agency. Additionally, this research introduces the Optimism-then-NoRegret framework, a novel contribution that integrates both our proposed methodologies and existing algorithms in the field.
@@
Figure 1: Samples required for different algorithms to reach the same average regret $(10^{-1})$ under a $50\times 50$ matrix with the self-play setting. (UCB-RM and OTS type are proposed algorithms. Moreover, IWE-Hedge corresponds to the Exp3 algorithm, and IWE-RM represents regret matching with bandit feedback. The UCB-Hedge algorithm was introduced by (Sessa et al., 2019).)
@@ 1 Introduction
Many real-world problems in economics (Fainmesser, 2012), sociology (Skyrms & Pemantle, 2009), transportation (Leblanc, 1975), politics (Ordeshook et al., 1986), signal processing (Song et al., 2011), and other fields (Fudenberg & Tirole, 1991) can be described as unknown games, where each player only observes their opponents’ actions and the noisy rewards associated with their own selected actions (referred to as bandit feedback). The goal of each player is to maximize their individual reward, and the only way to achieve this is to repeatedly play and learn the game structure from the corresponding observed rewards. The challenge in unknown games is how to efficiently learn from bandit feedback. Celebrated no-regret learning algorithms, such as Hedge (Freund & Schapire, 1997) and regret-matching (Hart & Mas-Colell, 2000), ensure sublinear regret guarantees under the full information setting, where the rewards of all actions at each round are observable. However, these algorithms cannot handle problems with bandit feedback, and faces the curse of multi-agency. This paper focuses on how to effectively utilize information on opponents’ actions and reward structures, which can substantially reduce budgets, as the example illustrated in Fig.1.
Table 1: Regret bounds comparison.
@@
Imagined Reward IWE UCB OTS [Ours]
No-Regret Update Hedge $\mathcal{O}\big{(}\sqrt{T\log\mathcal{A}}\big{)}$ $\mathcal{O}\big{(}\sqrt{T\mathcal{A}\log\mathcal{A}}\big{)}$ $\mathcal{O}\big{(}\sqrt{T\log\mathcal{A}}+\sqrt{\gamma_{T}\beta T}\big{)}$ $\mathcal{O}\big{(}\sqrt{T\log\mathcal{A}}+\sqrt{\gamma_{T}\beta T}\big{)}$
RM $\mathcal{O}\big{(}\sqrt{T\mathcal{A}}\big{)}$ $\mathcal{O}\big{(}T^{2/3}\mathcal{A}^{2/3}\big{)}$ [Ours] $\mathcal{O}\big{(}\sqrt{T\mathcal{A}}+\sqrt{\gamma_{T}\beta T}\big{)}$ [Ours] $\mathcal{O}\big{(}\sqrt{T\mathcal{A}}+\sqrt{\gamma_{T}\beta T}\big{)}$
Table 2: $\mathcal{A}$ and $\mathcal{B}$ are the sizes of action sets, $\gamma_{T}$ is the maximum information gain (Definition 4.3) and $\beta=\log(\mathcal{A}T)$.
@@ 1.1 Related Works
Adversarial bandit
In the full-information setting, multiplicative-weights (MW) algorithms such as Hedge (Freund & Schapire, 1997) can achieve optimal regret for adversarial bandit problems. Full information feedback requires perfect game knowledge and is unrealistic in many applications. In the more challenging bandit setting, Exp3 (Auer et al., 2002b) is a classical algorithm that utilizes the importance-weighted estimator for reward vector construction.
Game
@@
Previous literature ignores the fact that the agents are playing repeated games in which the reward structure could be exploited. Furthermore, to the application of our interest, the opponent’s actions can be observed by the agent. This scenario has received relatively little attention despite its numerous applications (O’Donoghue et al., 2021). Some papers consider a similar setting as ours. (O’Donoghue et al., 2021) seeks to compare the received reward to the Nash value and proposes variants of UCB and K-learning that converge to the Nash value. Our goal is to exploit the weakness of opponent and compete with hindsight’s best actions. With this performance metric, we seek to exploit the opponent’s strategy instead of only achieving Nash value. Similar to our setting, (Sessa et al., 2019) also uses adversarial regret and focuses on the single-player viewpoint. It utilizes Gaussian Process to exploit the correlations among different game outcomes and obtains a kernel-dependent regret bound with the factor $\gamma_{T}$ by an upper confidence bound (UCB) type algorithm. Thompson Sampling (TS) and other randomized exploration variants (Russo et al., 2018; Vaswani et al., 2020) are a strong counterpart of UCB-type algorithm in the reward structure-aware bandit literature. Moreover, (Zhang, 2022; Agarwal & Zhang, 2022) implements the optimism in Thompson sampling by introducing a optimism prior, which favors high reward models more aggressively than the standard Thompson sampling. However, this approach for optimistic TS is not computationally tractable. Another line of work is to tackle computation tractability issue of Thompson sampling (Li et al., 2022) for large-scale complex environment, but it remains the single-agent setups. However, no evidence shows these randomized exploration methods can work in our unknown game setting.
This work tries to understand how to efficiently learn to play unknown games. The main contributions are highlighted as follows:
  • Effective algorithm: We first demonstrate that vallina TS indeed fails in a specific class of unknown games. To overcome this, we introduce an optimistic variant of TS (referred as OTS for short) combined with appropriate full information adversarial bandit algorithms. We show that OTS can fix the divergence issue with the help of anti-concentration behavior of Gaussian distribution. Moreover, it significantly reduces experimental budgets in two practical problems. It decreases the budgets more than tenfold [^1] In Figure 5, OTS-Hedge achieves more than tenfold improvement comparing with IWE-hedge for reaching regret $10^{-2}$. compared to baseline algorithms in a two player radar sensing problem, and effectively mitigates the curse of multi-agency in traffic routing problem up to hundreds of decision-makers [^2] In Figure 6, OTS type and UCB type algorithms can effective handle hundreds of decision-makers..
  • Theoretical advantages: A general information-theoretic regret bound is provided and sublinear regret bounds of all proposed algorithms are established accordingly (Table 2). Our analysis highlights that using (i) the information of opponent’s actions and (ii) the underlying reward structure can help resolve the curse of multi-agency. For structured reward functions, our algorithms achieve regret bounds that depend logarithmically on the size of the action space. In contrast, algorithms that rely only on partial information suffer from curse of multi-agency.
  • Framework: An Optimism-then-NoRegret (OTN) learning framework for unknown games is also introduced. This framework encompasses various vanilla game algorithms as special cases, and several efficient algorithms can be developed under this framework, including upper confidence bound (UCB) based algorithms and TS-based algorithms. Notably, the proposed OTS-RM algorithm from this framework achieves the best performance in all experiments.
@@
This paper is organized as follows. Section 2 introduces the fundamental protocol and notations used in the repeated bandit game. Section 3 describes the proposed OTS type and UCB-RM algorithms, along with the corresponding OTN framework. Section 4 provides a regret analysis, while Section 5 details the experimental results. Section 6 concludes the paper.
@@ 2 Repeated Bandit Game
To simplify the presentation, we consider a two-player game scenario involving Alice and Bob. However, our results can be extended straightforwardly to multiplayer games by treating all other players as an abstract player.
@@
Consider a repeated game between Alice and her opponent Bob, where the action index sets for Alice and Bob are denoted by $\mathcal{A}={1,\ldots,\left|\mathcal{A}\right|}$ and $\mathcal{B}={1,\ldots,\left|\mathcal{B}\right|}$, respectively. [^3] We use the shorthand $\mathcal{A}$ to denote the cardinality $\left|\mathcal{A}\right|$. At each time $t=0,1,\ldots$, Alice selects an action $A_{t}\in\mathcal{A}$, and Bob simultaneously selects an action $B_{t}\in\mathcal{B}$. The mean reward for each action pair $(a,b)$ is $f_{\theta}(a,b)$, where $\theta\in\Theta$ is a model parameter unknown to the players. In the bandit feedback case, Alice only observes a noisy version of the mean reward associated with the selected action pair $(A_{t},B_{t})$
@@
where $\eta_{t}$ is an iid noise sequence. Under the full information setting, Alice can observe the whole mean rewards $\textbf{r}_{t}=(f_{\theta}(a,B_{t}))_{a\in\mathcal{A}}$ (details are presented in Appendix B). Alice’s experience up to time $t$ is encoded by the history $H_{t}=\left(A_{0},B_{0},Y_{1,A_{0},B_{0}},\ldots,A_{t-1},B_{t-1},Y_{t,A_{t-1},B_{t-1}}\right)$.
@@
An algorithm $\pi^{\operatorname{alg}}=(\pi_{t})_{t\in\mathbb{N}}$ employed by Alice is a sequence of deterministic functions, where each $\pi_{t}(H_{t})$ specifies a probability distribution over the action set $\mathcal{A}$ based on the history $H_{t}$. Alice’s action $A_{t}$ is sampled from the distribution $\pi_{t}$, i.e., $\mathbb{P}(A_{t}\in\cdot\>|\>\pi_{t})=\mathbb{P}(A_{t}\in\cdot\>|\>H_{t})=\pi_{t}(\cdot)$.
The above description of the bandit game encompasses various game forms based on the structure of the mean function $f_{\theta}(a,b)$. Several representative game forms such as the matrix game, linear game, and kernelized game are summarized in Table 4 (see Appendix B).
@@
[…]
$\sum_{t=0}^{T-1}\mathbb{E}\left[R_{t+1,A^{*},B_{t}}-R_{t+1,A_{t},B_{t}}\>|\>\theta\right],$ (1)
where the expectation is taken over the randomness in the actions $A_{t}$ and the rewards $R_{t+1,A_{t},B_{t}}$. However, this adversarial regret $\Re(T,\pi^{\operatorname{alg}},B_{0:T},\theta)$ is not a suitable regret metric under our game setting since it depends on the specific action sequence $B_{0:T}$ played by Bob. In this paper, we adopt the worse-case regret as the metric. An algorithm $\pi_{\operatorname{alg}}$ is considered No-Regret for Alice if, for any $B_{0:T}$, Alice suffers only sublinear regret, i.e.,
*$\Re^{*}(T,\pi^{\operatorname{alg}})=\sup_{B_{0:T}\in\mathcal{B}^{T}}\Re(T,\pi^{\operatorname{alg}},B_{0:T})=o(T),$
where $\theta$ is omitted for notation simplicity.
@@ 3.1 Review of Full Information Feedback
We will start by providing a brief overview of the full information feedback setting, in which Alice can observe the mean rewards $r_{t}(a)=f_{\theta}(a,B_{t})$ for all actions $a\in\mathcal{A}$ (details are presented in Appendix B). At time $t$, Alice picks action $A_{t}\sim P_{X_{t}}$ where $P_{X}(i)={X_{i}}/\sum_{i\in\mathcal{A}}X_{i}$. Full-information adversarial bandit algorithms, e.g., Hedge (Freund & Schapire, 1997) and Regret Matching (RM) (Hart & Mas-Colell, 2000), can be used to update $X_{t}$ to $X_{t+1}=g_{t}(X_{t},r_{t})$,
$\begin{cases}\textrm{Hedge:}\ &g_{t,a}(X_{t},r_{t})=X_{t,a}\exp\left(\eta_{t}r_{t}(a)\right),\\ \textrm{RM:}\ &g_{t,a}(X_{t},r_{t})=\max\left(0,\sum\limits_{s=0}^{t}r_{t}(a)-r_{t}(A_{s})\right),\end{cases}$
where $g_{t}(\cdot):\mathbb{R}_{+}^{\mathcal{A}}\times\mathbb{R}_{+}^{\mathcal{A}}\mapsto\mathbb{R}_{+}^{\mathcal{A}}$. The full information adversarial regret of an algorithm $\operatorname{adv}$ for a reward sequence $(r_{t})_{t}$ is defined as
$\Re_{\operatorname{full}}(T,{\operatorname{adv}},(r_{t})_{t})=\max_{a\in\mathcal{A}}\mathbb{E}\left[\sum_{t=0}^{T-1}r_{t}(a)-r_{t}(A_{t})\right].$ (2)
@@
$\Re_{\operatorname{full}}(T,{\operatorname{adv}})$ $=\max_{(r_{t})_{t}}\Re_{\operatorname{full}}(T,{\operatorname{adv}},(r_{t})_{t}).$ (3)
Since $r_{t}(a)=f_{\theta}(a,B_{t})$, the adversarial regret in Equation 1 can be reformulated as the full-information adversarial regret. For Hedge and RM (Freund & Schapire, 1997), their worst regret regrets can be bounded as $\Re_{\operatorname{full}}(T,{\operatorname{Hedge}})=\mathcal{O}(\sqrt{T\log\mathcal{A}})$ and $\Re_{\operatorname{full}}(T,{\operatorname{RM}})=\mathcal{O}(\sqrt{T\mathcal{A}})$.

### 25 Feb 2024 v2

@@ Abstract
This work tackles the complexities of multi-player scenarios in unknown games, where the primary challenge lies in navigating the uncertainty of the environment through bandit feedback alongside strategic decision-making. We introduce Thompson Sampling (TS)-based algorithms that exploit the information of opponents’ actions and reward structures, leading to a substantial reduction in experimental budgets—achieving over tenfold improvements compared to conventional approaches. Notably, our algorithms demonstrate that, given specific reward structures, the regret bound depends logarithmically on the total action space, significantly alleviating the curse of multi-player. Furthermore, we unveil the Optimism-then-NoRegret (OTN) framework, a pioneering methodology that seamlessly incorporates our advancements with established algorithms, showcasing its utility in practical scenarios such as traffic routing and radar sensing in the real world.
@@
Figure 1: Reduction on experimental budgets by our methods in terms of # samples required for various algorithms to reach the same level in average regret $(10^{-1})$ under the $50\times 50$ matrix game setups ( Section 5.1). IWE, UCB and OTS correspond to Importance Weighted Estimator, Upper Confidence Bound and Optimistic Thompson sampling. RM corresponds to Regret Matching. IWE-RM, UCB-RM, OTS-Hedge and OTS-RM are our proposed algorithms. The compared baselines include IWE-Hedge (known as Exp3  (Auer et al., 2002b) while we implement all Exp3 variants (Stoltz, 2005; Kocák et al., 2014; Lattimore & Szepesvári, 2020) and select the best as baseline.) and UCB-Hedge (known as GP-MW in (Sessa et al., 2019)).
@@ 1 Introduction
Many real-world problems in economics (Fainmesser, 2012), sociology (Skyrms & Pemantle, 2009), transportation (Leblanc, 1975), politics (Ordeshook et al., 1986), signal processing (Song et al., 2011), and other fields (Fudenberg & Tirole, 1991) can be described as unknown games, where each player only observes their opponents’ actions and the noisy rewards associated with their own selected actions (referred to as bandit feedback). The goal of each player is to maximize their individual reward, and the only way to achieve this goal is by repeatedly playing the game and learning its structure from the observed rewards. The challenge in unknown games is how to efficiently learn from bandit feedback. Celebrated no-regret learning algorithms, such as Hedge (Freund & Schapire, 1997) and regret-matching (Hart & Mas-Colell, 2000), ensure sublinear regret guarantees under the full information setting, where the rewards of all actions at each round are observable. However, these algorithms cannot handle problems with bandit feedback and face the curse of multi-player: the complexity scales exponentially as the number of players increases. We aim to develop efficient algorithms learning to play repeated unknown games by leveraging information on opponents’ actions and reward structures. We highlight the contributions:
  • Effective algorithms: The vanilla TS indeed fails in a specific class of unknown games as we demonstrated. To overcome this drawback, we introduce an optimistic variant of TS (a.k.a. OTS) combined with appropriate full information adversarial bandit algorithms. We show that OTS can fix the divergence issue with the help of Gaussian anti-concentration behavior. Empirically, our proposed methods reduce the experimental budgets more than an order of magnitude in two real-world problems - two-player radar sensing (Figure 5) and multi-player traffic routing problem (Figure 6) - and the random matrix game (Figures 1 and 3). Meanwhile, our proposed methodology effectively mitigates the curse of multi-player: OTS-type algorithms are capable of solving the traffic routing problem with hundreds of decision-makers (Figure 6), resulting in minimal congestion.
  • Theoretical advantages: A general information-theoretic regret bound is provided and sublinear regret bounds of all proposed algorithms are established accordingly (Table 1). Our analysis highlights that using (1) the information of opponent’s actions and (2) the underlying reward structure can help resolve the curse of multi-player. For structured reward functions, our algorithms achieve regret bounds that depend logarithmically on the size of the action space. In contrast, algorithms that rely only on bandit feedback suffer from the curse of multi-player.
  • Unified framework: An Optimism-then-NoRegret (OTN) learning framework for unknown games is also introduced. This framework encompasses various vanilla game algorithms as special cases, and several efficient algorithms can be developed under this framework, including upper confidence bound (UCB) based algorithms and TS-based algorithms. Notably, the proposed OTS-RM algorithm from this framework achieves the best performance in all experiments.
@@
[…]
No-Regret Update Hedge $\mathcal{O}\big{(}\sqrt{T\log\left|\mathcal{A}\right|}\big{)}$ $\mathcal{O}\big{(}\sqrt{T\mathcal{A}\log\left|\mathcal{A}\right|}\big{)}$ $\mathcal{O}\big{(}\sqrt{T\log\left|\mathcal{A}\right|}+\sqrt{\gamma_{T}\beta T}\big{)}$cellcolor[HTML]C1FEC0 $\mathcal{O}\big{(}\sqrt{T\log\left|\mathcal{A}\right|}+\sqrt{\gamma_{T}\beta T}\big{)}$
RM $\mathcal{O}\big{(}\sqrt{T\left|\mathcal{A}\right|}\big{)}$cellcolor[HTML]C1FEC0 $\mathcal{O}\big{(}T^{2/3}\left|\mathcal{A}\right|^{2/3}\big{)}$cellcolor[HTML]C1FEC0 $\mathcal{O}\big{(}\sqrt{T\left|\mathcal{A}\right|}+\sqrt{\gamma_{T}\beta T}\big{)}$cellcolor[HTML]C1FEC0 $\mathcal{O}\big{(}\sqrt{T\left|\mathcal{A}\right|}+\sqrt{\gamma_{T}\beta T}\big{)}$
Table 1: Comparison on the regret bounds. $\mathcal{A}$ is the action set of the agent and $\mathcal{B}$ is the action set of opponent. $\gamma_{T}$ is the maximum information gain (Definition 4.3) and $\beta=\log(\left|\mathcal{A}\right|T)$. The cells with green color are our results.
@@ 1.1 Related works
Adversarial Bandits.
In the full-information setting, multiplicative-weights (MW) algorithms such as Hedge (Freund & Schapire, 1997) achieve optimal regret for adversarial bandit problems. However, full information feedback, requiring perfect game knowledge, is unrealistic in many applications. In the challenging bandit feedback setting, the Exp3 algorithm (Auer et al., 2002b; Stoltz, 2005; Kocák et al., 2014; Neu, 2015; Lattimore & Szepesvári, 2020) is notable for utilizing an importance-weighted estimator to construct the reward vector.
Learning in Games.
@@
Structure & Opponent Awareness.
Prior literature often overlooks the potential to exploit the reward structure in repeated games and the observability of opponents’ actions. This oversight persists despite the scenario’s relevance to numerous applications (O’Donoghue et al., 2021; Sessa et al., 2019). (O’Donoghue et al., 2021) compares the received reward to the Nash value, proposing UCB and K-learning variants to minimize Nash regret. In contrast, our work aims to exploit the opponent’s strategy, introducing a focus on adversarial regret—a metric discussed further in Appendix A. This approach differs fundamentally from seeking only to achieve the Nash value. (Sessa et al., 2019) also prioritizes adversarial regret, employing a Gaussian Process to exploit correlations among game outcomes and achieve a kernel-dependent regret bound. This bound includes the factor $\gamma_{T}$, derived from an UCB-type algorithm. Thompson sampling (TS) and its variants (Russo et al., 2018; Vaswani et al., 2020) represent a strong alternative to UCB-type algorithms in the context of reward structure-aware bandit literature. Recent works by Zhang (2022); Agarwal & Zhang (2022) have explored the incorporation of optimism into TS through an optimistic prior, albeit facing computational tractability issues. Other research (Li et al., 2022, 2023, 2024) addresses these computational challenges in TS for large-scale, complex environments in single-agent setups. However, evidence supporting the effectiveness of these randomized exploration methods in our specific setting of unknown games remains limited.
Our proposed methodologies not only address the computational tractability issues inherent in optimistic TS but also introduce a novel perspective on learning in unknown game environments. By focusing on adversarial regret, we provide a more nuanced understanding of how players can strategically navigate these games to their advantage. This shift in focus from merely achieving Nash equilibrium to exploiting strategic opportunities represents a significant departure from traditional approaches.
@@
This paper is organized as follows: Section 2 introduces the fundamental protocols and notations used in the repeated bandit game. Section 3 describes the proposed OTS-type and UCB-RM algorithms, as well as the corresponding OTN framework. Section 4 provides an analysis of the regret associated with these methods. Section 5 details the experimental results, showcasing the effectiveness of our approaches. Finally, Section 6 concludes the paper.
@@ 2 Repeated bandit game
To simplify the exposition, we consider a two-player game scenario involving Alice and Bob. However, our results can be straightforwardly extended to multiplayer games by treating all other players as an abstract player.
@@
Consider a repeated game between Alice and her opponent, Bob, where the action index sets for Alice and Bob are denoted by $\mathcal{A}=\{1,\ldots,|\mathcal{A}|\}$ and =$\mathcal{B}=\{1,\ldots,|\mathcal{B}|\}$, respectively. [^1] We use the shorthand $|\mathcal{A}|$ to denote the cardinality of Alice’s action set, and similarly $|\mathcal{B}|$ for Bob’s. At each time $t=0,1,\ldots$, Alice selects an action $A_{t}\in\mathcal{A}$, and Bob simultaneously selects an action $B_{t}\in\mathcal{B}$. The mean reward for each action pair $(a,b)$ is $f_{\theta}(a,b)$, where $\theta\in\Theta$ is a model parameter unknown to the players. In the bandit feedback scenario, Alice only observes a noisy version of the mean reward associated with the selected action pair $(A_{t},B_{t})$:
@@
where $\eta_{t}$ is an i.i.d. noise sequence. Under the full information setting, Alice can observe the mean reward vector [^2] Details of the full information feedback protocol are presented in Appendix B $\mathbf{r}_{t}=(f_{\theta}(a,B_{t}))_{a\in\mathcal{A}}$ associated with each action $a\in\mathcal{A}$. Alice’s experience up to time $t$ is encoded by the history $H_{t}=\{(A_{0},B_{0},Y_{1,A_{0},B_{0}}),\ldots,(A_{t-1},B_{t-1},Y_{t,A_{t-1},B_{t-1}})\}$.
@@
An algorithm $\pi^{\text{alg}}=(\pi_{t})_{t\in\mathbb{N}}$ employed by Alice is a sequence of deterministic functions, where each $\pi_{t}(H_{t})$ specifies a probability distribution over the action set $\mathcal{A}$ based on the history $H_{t}$. Alice’s action $A_{t}$ is sampled from the distribution $\pi_{t}(H_{t})$, i.e., $\mathbb{P}(A_{t}\in\cdot|H_{t})=\pi_{t}(\cdot)$.
The above description of the bandit game encompasses various game forms based on the structure of the mean function $f_{\theta}(a,b)$. Several representative game forms, such as the matrix game, linear game, and kernelized game, are summarized in Table 3 (see Appendix B).
@@
Alice maintains a reward function $\mathcal{R}:\mathbb{R}\mapsto[0,1]$ that maps the observations to a bounded value, i.e., $R_{t+1,A_{t},B_{t}}=\mathcal{R}(Y_{t+1,A_{t},B_{t}})$. The objective for Alice is to maximize her expected reward $\sum_{t=0}^{T-1}\mathbb{E}[R_{t+1,A_{t},B_{t}}|\theta]$ over some duration $T$, irrespective of Bob’s fixed action sequence $B_{0:T}$. By treating Bob as the adversarial environment, the best action $A^{*}$ in hindsight is $A^{*}=\arg\max_{a\in\mathcal{A}}\sum_{t=0}^{T-1}\mathbb{E}[R_{t+1,a,B_{t}}|\theta]$, and the $T$-period adversarial regret $\mathcal{R}(T,\pi^{\text{alg}},B_{0:T},\theta)$ is defined by
$\sum_{t=0}^{T-1}\mathbb{E}[R_{t+1,A^{*},B_{t}}-R_{t+1,A_{t},B_{t}}|\theta],$ (1)
where the expectation is taken over the randomness in the actions $A_{t}$ and the rewards $R_{t+1,A_{t},B_{t}}$. However, this adversarial regret is not a suitable metric under our game setting since it depends on Bob’s specific action sequence $B_{0:T}$. We adopt the worst-case regret as the metric. An algorithm $\pi_{\text{alg}}$ is considered No-Regret for Alice if, for any $B_{0:T}$, Alice suffers only sublinear regret, i.e.,
*$\mathcal{R}^{*}(T,\pi^{\text{alg}})=\sup_{B_{0:T}\in\mathcal{B}^{T}}\mathcal{R}(T,\pi^{\text{alg}},B_{0:T})=o(T),$
omitting $\theta$ for simplicity of notation.
@@ 3.1 Review of Full Information Feedback
We start by providing a brief overview of the full information feedback setting, in which Alice can observe the mean rewards $r_{t}(a)=f_{\theta}(a,B_{t})$ for all actions $a\in\mathcal{A}$. At time $t$, Alice picks action $A_{t}\sim P_{X_{t}}$, where $P_{X}(i)=\frac{X_{i}}{\sum_{j\in\mathcal{A}}X_{j}}$. Full-information adversarial bandit algorithms, such as Hedge (Freund & Schapire, 1997) and Regret Matching (RM) (Hart & Mas-Colell, 2000), can be used to update $X_{t}$ to $X_{t+1}=g_{t}(X_{t},r_{t})$,
$\begin{cases}\text{Hedge:}&g_{t,a}(X_{t},r_{t})=X_{t,a}\exp\left(\eta_{t}r_{t}(a)\right),\\ \text{RM:}&g_{t,a}(X_{t},r_{t})=\max\left(0,\sum\limits_{s=0}^{t}r_{t}(a)-r_{t}(A_{s})\right),\end{cases}$
where $g_{t}(\cdot):\mathbb{R}_{+}^{\mathcal{A}}\times\mathbb{R}_{+}^{\mathcal{A}}\mapsto\mathbb{R}_{+}^{\mathcal{A}}$. The full information adversarial regret of an algorithm adv for a reward sequence $(r_{t})_{t}$ is defined as
$\mathcal{R}_{\text{full}}(T,\text{adv},(r_{t})_{t})=\max_{a\in\mathcal{A}}\mathbb{E}\left[\sum_{t=0}^{T-1}r_{t}(a)-r_{t}(A_{t})\right].$
@@
$\mathcal{R}_{\text{full}}(T,\text{adv})=\max_{(r_{t})_{t}}\mathcal{R}_{\text{full}}(T,\text{adv},(r_{t})_{t}).$ (2)
Since $r_{t}(a)=f_{\theta}(a,B_{t})$, the adversarial regret in Equation 1 can be reformulated as full-information adversarial regret. For Hedge and RM, their worst-case regrets can be bounded as $\mathcal{R}_{\text{full}}(T,\text{Hedge})=\mathcal{O}(\sqrt{T\log|\mathcal{A}|})$ and $\mathcal{R}_{\text{full}}(T,\text{RM})=\mathcal{O}(\sqrt{T|\mathcal{A}|})$.